# Robust Utility Optimization via a GAN Approach

### v1

@@ 1 Introduction
This approach offers more flexibility beyond fixed interval constraints for drift and diffusion terms. Nonetheless, the problem remains challenging to solve particularly in realistic market settings. First, analytic solutions even for these simple penalties are known only in standard cases, for log-preferences in friction-less markets (Guo et al., 2022). Second, classical numerical approaches quickly face limitations as the market dimension increases or trading costs are introduced.
@@
GANs have shown impressive results in various applications, including image and text generation, video synthesis, and drug discovery (Radford et al., 2016, Liu et al., 2020, Tripathi et al., 2022). Through robust approaches, GANs have progressed into mathematical finance as well (Cont et al., 2023). For robust utility optimization in a one-dimensional market, Guo et al. (2022) test a GAN approach for logarithmic utility functions with robust diffusion. In his simple setting, they find that their GAN approach recovers the (in this case known) analytic solution. A thorough application of GANs to multidimensional, realistic markets including trading costs in particular is yet outstanding. Moreover, the setup in Guo et al. (2022) contains some unrealistic assumptions on NN architectures that essentially mean the antagonistic market is fully aware of the generator’s trades and vice versa. Hence there is a lot of room for an in-depth study of GAN technologies in robust portfolio optimization. For simplicity, we speak of the GAN approach or GAN strategies even if we fix the market dynamics and only (non-robustly) train NN approximations of optimal investment strategies.
One major advantage of the GAN approach to (robust) utility optimization is that it is universally applicable across choices of utility functions, market structures, costs and other constraints. In particular, this allows us to apply our method to the problem of optimal investment under transaction costs (Constantinides, 1979, Davis and Norman, 1990, De Lataillade et al., 2012, Tunc et al., 2013, Martin, 2012, 2016, Muhle-Karbe et al., 2017, Belak et al., 2022, Sun et al., ), and compare our soultions to well-known asymptotically optimal (for vanishing trading costs), non-robust strategies that are characterized by no-trade regions around the optimal friction-less Merton strategy (Merton, 1969).
@@ 1.1 Overview of our Contribution
  • We further analyze trained strategies for markets with trading costs and general power utilities, where no analytical solution is known.
@@
  • Without analytic solutions the results in these more realistic settings require proper evaluation metrics. We propose several metrics for evaluating the learned investment strategies based on a pool of potential market models.
@@
This paper is structured as follows. We first introduce our continuous-time financial market model and the penalized robust utility optimization problem including transaction costs in Section 2. Then, in Section 3, we outline our GAN-based algorithm for robust utility optimization in discretized time. In Section 4, we discuss the effectiveness of the GAN-based approach in various experiments: In Section 4.1, we test the approach in (friction-less) settings with known analytic solutions, in Section 4.2 we test it in markets with proportional trading costs, and Section 4.3 contains experiments on robust utility optimization with path-wise penalties. Moreover, we further study the effects of small transaction costs in robust and non-robust utility optimization in Section 4.4. Finally, we conclude our findings in Section 5.
@@ 2 Preliminaries
[…]
@@
for a given utility function $u$ and where $\mathcal{P}$ is a family of possible underlying probabilistic models $\mathbb{P}$. The family $\mathcal{P}$ represents epistemic uncertainty. It contains measures $\mathbb{P}$ that generate plausible laws of the underlying assets $S$. In particular, we assume that *$\mathbb{P}^{*}\in\mathcal{P}$.
@@
The number of measures constructed in that way is vast, and further restrictions on $\mathcal{P}$ are in order to arrive at tractable solutions. Perhaps the most realistic approach to set up potential market measures is to add constraints on $\mathbb{P}\in\mathcal{P}$ that tie measures $\mathbb{P}$ to observed market behavior We incorporate such constraints on $\mathcal{P}$ by adding penalty terms to the objective of Equation 2. More formally, for a penalization functional $R$, let the constrained version of Equation 2 be given as
@@
[…]
@@
However, penalties of the type of Equation 5 face the criticism that reference values, in particular for the drift component $\mu$, can not be estimated well. By extension, these quantities can not be inferred from observing the market, and searching a pool of models tied to these characteristics is not in line with the initial aim. Consequently, the following question arises. What are adequate market measures to search for in robust portfolio optimization and how do we need to choose the penalties to make this search plausible?
This question is deeply rooted in the field of stochastic portfolio theory (SPT) pioneered by Fernholz and Karatzas (2009). The central aim of SPT is to develop methods for portfolio optimization that only depend on observable quantities. In SPT, observable quantities are prices, but also properties of paths that can be estimated well. For instance, this includes covariances along an observed trajectory, or some invariant laws observable in the market. However, drift coefficients are not considered observable. Thus, the SPT approach differs from the one of Equation 5 that ties both drift and diffusion terms to certain reference values.
A fairly general way to incorporate observable quantities into the robust utility maximization problem of Equation 2 is to include path-wise constraints in $\mathcal{P}$. These constraints can be constructed to relate a potential market measure $\mathbb{P}$ for $S$ to an observed trajectory $S^{\mathrm{obs}}$. The idea is to consider any functional $l:C([0,t],\mathbb{R}^{d})\to\mathbb{R}$ with $t\in[0,T]$ and set up a constraint to require that $l(S)$ is close to $l(S^{\mathrm{obs}})$. Here, closeness can for instance be defined via an $L_{2}$-norm or in $\mathbb{P}$-expectation. In terms of soft constraints, these path-wise penalties can be incorporated into Equation 4 via
[…]
@@
In this paper, we investigate the robust utility optimization problem with path-wise constraints building on the functionals of quadratic covariation up to terminal time $T$, i.e., $l_{\text{QCV}}(S)=[S]_{T}$ (with penalty $R_{1}$), and relative return over the entire time horizon $l_{\text{RR}}(S)=\frac{S_{T}}{S_{0}}$ (with penalty $R_{2}$). In particular, the latter boils down to computing the average relative return $l_{\text{ARR}}(S)=\mathbb{E}_{\mathbb{P}}\left[l_{\text{RR}}(S)\right]=\mathbb{E}_{\mathbb{P}}\left[\frac{S_{T}}{S_{0}}\right]$ (see also Section 4.3.2).
@@ 2.1 Related Work
An instance of the “penalization approach” of Equation 5 is proposed in Guo et al. (2022). In this framework, analytical solutions are known for some special cases. In particular, Guo et al. (2022) derive exact solutions to the robust optimization problem for logarithmic utilities and $\lambda_{2}=0$. Moreover in this setting, Guo et al. (2022) are the first to demonstrate that a deep learning method based on GANs can recover the robust optimal strategy in a one-dimensional market. Nonetheless, solving a robust utility maximization problem in general markets including frictions, for general utility functions and constraints remains a challenging task. Neither analytical solutions are known, nor do numerical solutions exist so far. Moreover, to the best of our knowledge, the robust utility optimization problem with path-wise constraints of the form of Equation 6 or Equation 7 has not yet been analyzed numerically.
@@ 3 Deep Robust Utility Optimization
Deep learning has become increasingly popular in mathematical finance for a variety of tasks that are not easily solvable with conventional methods. In the context of robust utility optimization, such an intricate task is finding trading strategies that are robust to adversarial market movements in realistic markets with trading costs. In this section, we first introduce the (realistic) discrete-time market setting for approximating solutions to the penalized robust utility optimization problem Equation 4 in discrete time. Secondly, we propose a GAN-based algorithm for approximating solutions to this problem. This includes approximations of both adversarial market measures and robust trading strategies.
@@ 3.1 Discrete-time Market Setting
For numerically solving the penalized robust utility optimization problem Equation 4, we discretize the time interval $[0,T]$ into $N\in\mathbb{N}$ time steps $0=t_{0}<\ldots<t_{N}=T$, with step sizes $\Delta t_{n}:=t_{n}-t_{n-1}$. Let the discrete-time dynamics of assets $S=(S_{t_{n}})_{n=0,\ldots,N}$ be given by the Euler-Maruyama scheme as
@@
Here, we include proportional transaction costs $c_{\text{prop}}$ and additive base costs $c_{\text{base}}$ per trade, with asset dynamics as in Equation 8. Moreover, trading strategies are parametrized as portfolio weights $\pi=(\pi_{t_{n}})_{n=1,\ldots,N}$, where $\pi_{t_{n}}$ are $\mathcal{F}_{n}$-measurable functions taking values in $\mathbb{R}^{d}$.
@@ 3.2 GAN-based Algorithm
The discrete-time market setting outlined in Section 3.1 serves as basis for numerical approaches to approximate saddle point solutions of problem (4). In this setting, we look for robust trading strategies $\pi=(\pi_{t_{n}})_{n=1,\ldots,N}$ and adversarial market measures $\mathbb{P}_{\mu,\sigma}$ parametrized by $(\mu,\sigma)=(\mu_{n},\sigma_{n})_{n=0,\ldots,N-1}$, i.e., we search for saddle point solutions to the discrete-time min-max problem
$\max_{\pi}\min_{\mathbb{P}_{\mu,\sigma}}\mathbb{E}_{\mathbb{P}_{\mu,\sigma}}\left[u(X_{T}^{\pi,\mu,\sigma})\right]+R(\mathbb{P}_{\mu,\sigma}).$ (10)
@@
7:     gen loss $=-\left(\frac{1}{B}\sum u(x_{T})\right)-\lambda R_{T}$
8:     disc loss =absent=-= -gen loss
9:     if modulo(niter, 2) ==0==0= = 0 then
[…]
@@
[…]
@@
2:  asset_paths =[]absent=[]= [ ]
3:  market_measures =[]absent=[]= [ ]
4:  for $t=0$ to $T-1$ do
@@
7:     $s_{t+1},x_{t+1}=$ step$(s_{t},x_{t},a_{t},\mu_{t},\sigma_{t},D_{B})$
8:     asset_paths $=$ asset_paths $\cup\,s_{t+1}$collect asset values}
9:     market_measures $=$ market_measures $\cup\,\mu_{t},\sigma_{t}$collect market measures}
@@
11:  $R_{T}=$ estimate_path_functional(asset_paths, market_measures)
@@
An outline of a GAN algorithm for robust utility optimization is given in Algorithms 1 and 2. Following initializations of generator and discriminator networks in lines 2-3 of Algorithm 1, these networks are trained on a pre-generated training set of Brownian increments $D$ in lines 4-13 of Algorithm 1. For each iteration of the training, a batch of Brownian increments is selected from the training data in line 5 of Algorithm 1. Then, in a forward pass through generator and discriminator networks in code line 6, the terminal portfolio value as well as an estimate for the penalty term $R$ are determined for all increments in the current batch $D_{B}$. Algorithm 2 describes the forward pass in which paths for asset values and market measures are collected by consecutive evaluations of generator and discriminator networks. We remark that for our experiments we actually use the special form
@@
of the asset dynamics (1) (where $\odot$ is the Hadamard product), as can be seen in line 6 of Algorithm 2. The step function in line 7 of Algorithm 2 is given by (8). Having obtained realizations of terminal portfolio values and estimates of the penalty term based on the training increments $D_{B}$ in the forward pass, losses for generator and discriminator networks are calculated in lines 7-8 of Algorithm 1 as Monte Carlo estimates of the (negative) objective Equation 11. Finally, the NNs are updated in code lines 9-13 alternately over iterations. Naturally, there are various specifics to be set for these updates such as optimization algorithms and regularization methods. We refer to Section 4 for details on these choices in our experiments.
Moreover, throughout all experiments in Section 4, we put particular focus on uncovering potential path-dependencies in the optimized robust trading strategies (and the optimized characteristics of the adversarial market measure). To that end, we contrast recurrent and feed forward NN architectures for both discriminator and generator networks $(\mu^{\phi},\sigma^{\phi})$ and $\pi^{\theta}$, respectively.

### v2

@@ 1 Introduction
This approach offers more flexibility beyond fixed interval constraints for drift and diffusion terms. Nonetheless, the problem remains challenging to solve particularly in realistic market settings. First, explicit solutions even for these simple penalties are known only in standard cases, for log-preferences in friction-less markets (Guo et al., 2022). Second, classical numerical approaches quickly face limitations as the market dimension increases or trading costs are introduced.
@@
GANs have shown impressive results in various applications, including image and text generation, video synthesis, and drug discovery (Radford et al., 2016, Liu et al., 2020, Tripathi et al., 2022). Through robust approaches, GANs have progressed into mathematical finance as well (Cont et al., 2023). For robust utility optimization in a one-dimensional market, Guo et al. (2022) test a GAN approach for logarithmic utility functions with robust diffusion. In his simple setting, they find that their GAN approach recovers the (in this case known) explicit solution. A thorough application of GANs to multidimensional, realistic markets including trading costs in particular is yet outstanding. Moreover, the setup in Guo et al. (2022) contains some unrealistic assumptions on NN architectures that essentially mean the antagonistic market is fully aware of the generator’s trades and vice versa. Hence there is a lot of room for an in-depth study of GAN technologies in robust portfolio optimization. For simplicity, we speak of the GAN approach or GAN strategies even if we fix the market dynamics and only (non-robustly) train NN approximations of optimal investment strategies.
One major advantage of the GAN approach to (robust) utility optimization is that it is universally applicable across choices of utility functions, market structures, costs and other constraints. In particular, this allows us to apply our method to the problem of optimal investment under transaction costs (Constantinides, 1979, Davis and Norman, 1990, De Lataillade et al., 2012, Tunc et al., 2013, Martin, 2012, 2016, Muhle-Karbe et al., 2017, Belak et al., 2022, Sun et al., 2019), and compare our soultions to well-known asymptotically optimal (for vanishing trading costs), non-robust strategies that are characterized by no-trade regions around the optimal friction-less Merton strategy (Merton, 1969).
@@ 1.1 Overview of our Contribution
  • We further analyze trained strategies for markets with trading costs and general power utilities, where no explicit solution is known.
@@
  • Without explicit solutions the results in these more realistic settings require proper evaluation metrics. We propose several metrics for evaluating the learned investment strategies based on a pool of potential market models.
@@
This paper is structured as follows. We first introduce our continuous-time financial market model and the penalized robust utility optimization problem including transaction costs in Section˜2. Then, in Section˜3, we outline our GAN-based algorithm for robust utility optimization in discretized time. In Section˜4, we discuss the effectiveness of the GAN-based approach in various experiments: In Section˜4.1, we test the approach in (friction-less) settings with known explicit solutions, in Section˜4.2 we test it in markets with proportional trading costs, and Section˜4.3 contains experiments on robust utility optimization with path-wise penalties. Moreover, we further study the effects of small transaction costs in robust and non-robust utility optimization in Section˜4.4. Finally, we conclude our findings in Section˜6.
@@ 2 Preliminaries
$\Pi:=\left\{\pi=(\pi_{t})_{t\in[0,T]}:\pi_{t}\in\mathbb{R}^{d},\pi\text{ is }\mathbb{F}\text{-predictable}\right\}.$
@@
for a given utility function $u$ and where $\mathcal{P}$ is a family of possible underlying probabilistic models $\mathbb{P}$. The family $\mathcal{P}$ represents epistemic uncertainty. It contains measures $\mathbb{P}$ that generate plausible laws of the underlying assets $S$. In particular, we assume that $\mathbb{P}^{*}\in\mathcal{P}$.
@@
The number of measures constructed in that way is vast, and further restrictions on $\mathcal{P}$ are in order to arrive at tractable solutions. Perhaps the most realistic approach to set up potential market measures is to add constraints on $\mathbb{P}\in\mathcal{P}$ that tie measures $\mathbb{P}$ to observed market behavior We incorporate such constraints on $\mathcal{P}$ by adding penalty terms to the objective of Equation˜2. More formally, for a penalization functional $R$, let the constrained version of Equation˜2 be given as
@@
$R_{\lambda_{1},\lambda_{2}}(\mathbb{P}_{\mu,\sigma})=\mathbb{E}_{\mathbb{P}_{\mu,\sigma}}\left[\lambda_{1}\int_{0}^{T}\lVert\sigma_{t}\sigma_{t}^{\top}-\tilde{\sigma}\tilde{\sigma}^{\top}\rVert_{F}^{2}\,dt+\lambda_{2}\int_{0}^{T}\lVert\mu_{t}-\tilde{\mu}\rVert_{2}^{2}\,dt\right],$ (5)
@@
However, penalties of the type of Equation˜5 face the criticism that reference values, in particular for the drift component $\mu$, can not be estimated well. By extension, these quantities can not be inferred from observing the market, and searching a pool of models tied to these characteristics is not in line with the initial aim. Consequently, the following question arises. What are adequate market measures to search for in robust portfolio optimization and how do we need to choose the penalties to make this search plausible?
This question is deeply rooted in the field of stochastic portfolio theory (SPT) pioneered by Fernholz and Karatzas (2009). The central aim of SPT is to develop methods for portfolio optimization that only depend on observable quantities. In SPT, observable quantities are prices, but also properties of paths that can be estimated well. For instance, this includes covariances along an observed trajectory, or some invariant laws observable in the market. However, drift coefficients are not considered observable. Thus, the SPT approach differs from the one of Equation˜5 that ties both drift and diffusion terms to certain reference values.
A fairly general way to incorporate observable quantities into the robust utility maximization problem of Equation˜2 is to include path-wise constraints in $\mathcal{P}$. These constraints can be constructed to relate a potential market measure $\mathbb{P}$ for $S$ to an observed trajectory $S^{\mathrm{obs}}$. The idea is to consider any functional $l:C([0,t],\mathbb{R}^{d})\to\mathbb{R}$ with $t\in[0,T]$ and set up a constraint to require that $l(S)$ is close to $l(S^{\mathrm{obs}})$. Here, closeness can for instance be defined via an $L_{2}$-norm or in $\mathbb{P}$-expectation. In terms of soft constraints, these path-wise penalties can be incorporated into Equation˜4 via
$R_{1}(\mathbb{P})$ $=\lambda_{1}\int_{\Omega}\left\|{l}(S)-{l}(S^{\mathrm{obs}})\right\|_{2}^{2}\,d\mathbb{P},\text{ or}$ (6)
@@
In this paper, we investigate the robust utility optimization problem with path-wise constraints building on the functionals of quadratic covariation up to terminal time $T$, i.e., $l_{\text{QCV}}(S)=[S]_{T}$ (with penalty $R_{1}$), and relative return over the entire time horizon $l_{\text{RR}}(S)=\frac{S_{T}}{S_{0}}$ (with penalty $R_{2}$). In particular, the latter boils down to computing the average relative return $l_{\text{ARR}}(S)=\mathbb{E}_{\mathbb{P}}\left[l_{\text{RR}}(S)\right]=\mathbb{E}_{\mathbb{P}}\left[\frac{S_{T}}{S_{0}}\right]$ (see also Section˜4.3.2).
Remark 2.1.
If the penalty is of the type (5) where the drift is fixed and not optimized and $u$ is a continuous, increasing and concave utility function on $\mathbb{R}$, then it was shown in Guo et al. (2022, Proposition 2) that the $\sup\inf$ and $\inf\sup$ formulation of (4) coincide.
@@ 2.1 Related Work
An instance of the “penalization approach” of Equation˜5 is proposed in Guo et al. (2022). In this framework, explicit solutions are known for some special cases. In particular, Guo et al. (2022) derive exact solutions to the robust optimization problem for logarithmic utilities and $\lambda_{2}=0$. Moreover in this setting, Guo et al. (2022) are the first to demonstrate that a deep learning method based on GANs can recover the robust optimal strategy in a one-dimensional market. Nonetheless, solving a robust utility maximization problem in general markets including frictions, for general utility functions and constraints remains a challenging task. Neither explicit solutions are known, nor do numerical solutions exist so far. Moreover, to the best of our knowledge, the robust utility optimization problem with path-wise constraints of the form of Equation˜6 or Equation˜7 has not yet been analyzed numerically.
@@
Limmer and Horvath (2023) use a GAN framework applied to a penalized version of the robust hedging problem. In particular, similarly to our work, Limmer and Horvath (2023) use two neural networks competing in a penalized zero-sum game, where one models the market and the other one the trading (or hedging) strategy. While our strategy’s objective is to maximize the robust expected utility of the terminal wealth, Limmer and Horvath (2023) try to minimize the robust distance (measured via a utility functional) between the payoff of a derivative and the wealth achieved with the hedging strategy at terminal time. Robust hedging is also considered by Wu and Jaimungal (2023) under a similar robustification setting as in Jaimungal et al. (2022).
@@ 3 Deep Robust Utility Optimization
Deep learning has become increasingly popular in mathematical finance for a variety of tasks that are not easily solvable with conventional methods. In the context of robust utility optimization, such an intricate task is finding trading strategies that are robust to adversarial market movements in realistic markets with trading costs. In this section, we first introduce the (realistic) discrete-time market setting for approximating solutions to the penalized robust utility optimization problem Equation˜4 in discrete time. Secondly, we propose a GAN-based algorithm for approximating solutions to this problem. This includes approximations of both adversarial market measures and robust trading strategies.
@@ 3.1 Discrete-time Market Setting
For numerically solving the penalized robust utility optimization problem Equation˜4, we discretize the time interval $[0,T]$ into $N\in\mathbb{N}$ time steps $0=t_{0}<\ldots<t_{N}=T$, with step sizes $\Delta t_{n}:=t_{n}-t_{n-1}$. Let the discrete-time dynamics of assets $S=(S_{t_{n}})_{n=0,\ldots,N}$ be given by the Euler-Maruyama scheme as
@@
Here, we include proportional transaction costs $c_{\text{prop}}$ and additive base costs $c_{\text{base}}$ per trade, with asset dynamics as in Equation˜8. Moreover, trading strategies are parametrized as portfolio weights $\pi=(\pi_{t_{n}})_{n=1,\ldots,N}$, where $\pi_{t_{n}}$ are $\mathcal{F}_{n}$-measurable functions taking values in $\mathbb{R}^{d}$.
@@ 3.2 GAN-based Algorithm
The discrete-time market setting outlined in Section˜3.1 serves as basis for numerical approaches to approximate saddle point solutions of problem (4). In this setting, we look for robust trading strategies $\pi=(\pi_{t_{n}})_{n=1,\ldots,N}$ and adversarial market measures $\mathbb{P}_{\mu,\sigma}$ parametrized by $(\mu,\sigma)=(\mu_{n},\sigma_{n})_{n=0,\ldots,N-1}$, i.e., we search for saddle point solutions to the discrete-time min-max problem
$\sup_{\pi}\inf_{\mathbb{P}_{\mu,\sigma}}\mathbb{E}_{\mathbb{P}_{\mu,\sigma}}\left[u(X_{T}^{\pi,\mu,\sigma})\right]+R(\mathbb{P}_{\mu,\sigma}).$ (10)
@@
While the order of $\sup$ and $\inf$ in (10) can be switched in special cases (˜2.1), this is not necessarily true in general. From the perspective of robust utility optimisation, we are interested in the problem as stated in (10), i.e., in finding a worst-case optimal trading strategy. It is important to note that, in general, the GAN learns neither the $\sup\inf$ nor the $\inf\sup$ [^2] For the GAN to learn e.g. the $\sup\inf$ of Equation 10, the discriminator would need to be trained until it converged to the global minimum after each generator step; and vice versa for learning $\inf\sup$. It can be tried to achieve approximations of these by an imbalance between the number of steps for generator and discriminator.. Instead it tries to find a regularized [^3] More precisely, the problem is regularized through the parameterisation of the player’s strategies via neural networks and at a stationary point (in the parameter space) the derivative of the objective function with respect to the trainable parameters of each player is 0. stationary point, which can be interpreted as a local Nash equilibrium. However, it is well known that the training of GANs is unstable and usually not even a stationary point is found. But at the same time, the GAN is likely to heavily explore the space of trading strategies. We make use of this by early stopping criteria based on validation metrics approximating the worst-case expected utility. Hence, the trained, early-stopped strategy should approximate the worst-case optimal strategy.
@@
7:  gen loss $\left(\frac{1}{B}\sum u(x_{T})\right)-\lambda R_{T}$
8:  disc loss ==-gen loss
9:  if modulo(niter, 2) ==0==0 then
10:   $\pi^{\theta_{k+1}}=\texttt{train\textunderscore NN}$($\theta_{k}$, gen loss)
@@
12:   $\mu^{\phi_{k+1}},\sigma^{\phi_{k+1}}=\texttt{train\textunderscore NN}$($\phi_{k}$, disc loss)
@@
2: asset_paths =[]=[]
3: market_measures =[]=[]
4:for $t=0$ to T1T-1 do
@@
[…]
8:  asset_paths == asset_paths $\cup\,s_{t+1}$ {collect asset values}
9:  market_measures == market_measures $\cup\,\mu_{t},\sigma_{t}$ {collect market measures}
@@
[…]
@@
An outline of a GAN algorithm (without early stopping criterion) for robust utility optimization is given in Algorithms˜1 and 2. Following initializations of generator and discriminator networks in lines 2-3 of Algorithm˜1, these networks are trained on a pre-generated training set of Brownian increments $D$ in lines 4-13 of Algorithm˜1. For each iteration of the training, a batch of Brownian increments is selected from the training data in line 5 of Algorithm˜1. Then, in a forward pass through generator and discriminator networks in code line 6, the terminal portfolio value as well as an estimate for the penalty term $R$ are determined for all increments in the current batch $D_{B}$. Algorithm˜2 describes the forward pass in which paths for asset values and market measures are collected by consecutive evaluations of generator and discriminator networks. We remark that for our experiments we actually use the special form
@@
of the asset dynamics (1) (where $\odot$ is the Hadamard product), as can be seen in line 6 of Algorithm˜2. The step function in line 7 of Algorithm˜2 is given by (8). Having obtained realizations of terminal portfolio values and estimates of the penalty term based on the training increments $D_{B}$ in the forward pass, losses for generator and discriminator networks are calculated in lines 7-8 of Algorithm˜1 as Monte Carlo estimates of the (negative) objective Equation˜11. Finally, the NNs are updated in code lines 9-13 alternately over iterations. Naturally, there are various specifics to be set for these updates such as optimization algorithms and regularization methods. We refer to Section˜4 for details on these choices in our experiments.
Moreover, throughout all experiments in Section˜4, we put particular focus on uncovering potential path-dependencies in the optimized robust trading strategies (and the optimized characteristics of the adversarial market measure). To that end, we contrast recurrent and feed forward NN architectures for both discriminator and generator networks $(\mu^{\phi},\sigma^{\phi})$ and $\pi^{\theta}$, respectively.